OpenStudy (anonymous):

The average (arithmetic mean) of 5 numbers is 13. If the average of 4 of these numbers is 10, what is the fifth number?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

Let x be the sum of the first four numbers and y the fifth number. Solve the following for x and y:\[\left\{\frac{x + y}{5}=13,\frac{x}{4}=10\right\} \]\[\{x=40,y=25\} \]